if sec4A=cosec(A-20); where 4A is an acute angle find the value of A

Grade:10

1 Answers

Taushif alam
6 years ago
We have => sec4A =cosec(A-20)According to trigonometrical identities[ sec theta=cosec( 90-theta)]SoSec4A=cosec(A-20)Cosec(90-4A) = cosec(A-20)(90-4A)=(A-20)90-4A = A-20110=5AA=110/5A=22Hence the value of A = 22.